An electric circuit is formed of :

_a generator delivering a constant voltage Ugenerator=24 volt

_Resistor D of resistance R=48 ohm

_Rheostat (D') of adjustable resistance Rh


1_Show applying law of addition of voltages that the current is given by the expression:

I=Ugenerator/R+Rh


2_We vary the value of Rh between two limiting values 0 ohm and 120 ohm.Calculate the value I1 of I for Rh=0 ohm.Then calculate the value I2 of I for Rh=120 ohm.


3_Deduce the role of a rheostat in a series connection.


4_Let P be the power dissipated by R.Give the expression of P in terms of R and I.Calculate the value P1 of P for I=I1 and the value of P2 of P for I=I2


5_The maximum power that the resistor D can stand is 5 watt. Show that resistor D can be damaged only in one of the two limiting values of Rh.




1
Expert's answer
2021-12-22T14:14:08-0500

Let "I" = the current in the circuit

Part 1


pd across "D= IR"

pd across "D'=IR_h"

"IR+IR_h= U_{generator}\\I= \frac{U_{generator}}{R+R_h}"

Part 2


"I_1=\frac{24}{48+0}=0.5A"


"I_2= \frac{24}{48+120}=0.1429A"


Part 3


It regulates the current flowing in a circuit


Part 4


"P= I^2R"


"P_1= 0.5^2\u00d748=12 W"


"P_2=O.1429^2\u00d748=0.9802W"


Part 5


When "R_h=0\Omega\>\>I_1=0.5A\>\>and"

"P=12W"


Since 12W>5W, the resistor D will be damaged by lower limit of "R_h"


When "R_h=120\Omega\,\,I_2=0.1429A"

and "P=0.9802W"


0.9802W<5W

The resistor D will not be damaged by upper limit of "R_h"

Therefore, the resistor D can be damaged only in one of the two limiting values of the "R_h"